OpenStudy (anonymous):

Solve for x. 3^2x = 27

jimthompson5910 (jim_thompson5910):

3^2x = 27 3^2x = 3^3 Since the bases are equal (to 3), the exponents must be equal. So 2x = 3 Solve this for x to find your answer.

OpenStudy (anonymous):

So would it be 3/2?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

It says its wrong :(

jimthompson5910 (jim_thompson5910):

then it probably wants a decimal answer

OpenStudy (anonymous):

3 exp 2= 9 9x=27 27/9=3 x=3

jimthompson5910 (jim_thompson5910):

so try 3/2 = 1.5

OpenStudy (anonymous):

That was it :) Thanks!

jimthompson5910 (jim_thompson5910):

ah that's great

jimthompson5910 (jim_thompson5910):

np